VanRee Recycling Inc. has the following income statement for this year: Income Statement 750,000 (218,000) (190,000) 342,000 (150,000) 192,000 (57,600) 134,400 Sales Costs Depreciation EBIT Interest expense EBT Taxes Net Income Dividends 100,800 Additions to Retained Earnings 33,600 VanRee expects sales to rise by 10 percent next year. Costs are expected to increase spontaneously with sales Depreciation and interest expense will remain constant. The tax rate will also remain constant. VanRee will continue to pay out 75 percent of their net income as dividends Prepare next years income statement. What will their additions to retained earnings be? Round your answer to the nearest whole number and no commas Answer:
Answer #1

Answer:

Tax rate=57600/192000=30%

Sales=1.1*previous year question =1.1*750000=$825000

Cost=1.1*previous year question =1.1*218000=$239800

Income Statement
Sales 825000
Costs 239800
Depreciation 190000
EBIT =S-C-D 395200
Interest Expenses  I 150000
EBT =EBIT - I 245200
Taxes T=30%*EBT 73560
Net Income N= EBT-T 171640
Dividend =75%*N 128730
Addition to retained earning=25%*N 42910

Addition to retained earning =$42910